# Entanglement classification and non-k-separability certification via Greenberger-Horne-Zeilinger-class fidelity

### v1

@@ III.2 Four-qubit entanglement classification
For each of the nine classes, we maximize the correlator $\mathcal{Q}$ by picking the maximally entangled representatives of each group (Appendix D and Ref. [Vintskevich_2023]). As shown in Table 3, the highest value of the correlator is reached by the GHZ state ($|G\rangle$ group), and $|E_{1}\rangle$ and $|E_{2}\rangle$ groups, while the smallest $\mathcal{Q}$ is for group $|E_{6}\rangle$.
@@
Table 3: Value of the correlator $\mathcal{Q}$ for maximally entangled four-qubit states in nine entanglement classes (see App. D). The first column corresponds to the value of the correlator of the full state, while the latter four show $\mathcal{Q}$ after discarding (tracing out) one qubit. Null values corresponds to negative lack of entanglement. Observe that the state exhibiting the highest entanglement with respect to four-qubit correlator ($|G\rangle$) does not possess any three-qubit entanglement. Conversely, the states with lower four-qubit entanglement preserve high values of the three-qubit correlator. This is in agreement with similar observations concerning distribution of entanglement among different divisions [Verstraete_2002].
@@
Here, we consider only the qubit case, $d=2$. In the case of $N=2,3$ qubits, the AME is the corresponding GHZ state. Surprisingly, there are no four-qubit AME states [Higuchi_2000], and there are no AME states for $N>6$ [Scott_2004, Huber_2017]; but they do exist for $N=5,6$ [Facchi_2008, Goyeneche_2015], taking the form of
@@
where the values of binary (±$\pm 1$) coefficients $\alpha^{(5)}_{i}$ and $\alpha^{(6)}_{i}$ are given in Appendix E. For ${\rm AME}(5,2)$ the correlator takes $\mathcal{Q}=1.7$ and for ${\rm AME}(6,2)$ it takes value $\mathcal{Q}=2.4$. These values are about half of maximal values for $N=5$ and $N=6$ qubits ($\mathcal{Q}=4$ and $\mathcal{Q}=5$ respectively, see Table 1.).
@@ Acknowledgement
We thank Guillem Müller-Rigat, Anna Sanpera, Geza Toth, and Karol Życzkowski for reading the manuscript and their useful comments.
ICFO group acknowledges support from: Europea Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA MAQS PCI2019-111828-2); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733); Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID 100010434): CF/BQ/PR23/11980043.
@@ Appendix A Convexity of correlator $\mathcal{Q}$
$\begin{split}2\log_{4}\bigg{(}\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}&|\langle\tilde{0}|^{\otimes N}(\lambda\hat{\varrho}_{1}+(1-\lambda)\hat{\varrho}_{2})|\tilde{1}\rangle^{\otimes N}|\bigg{)}\leq\\ \leq 2\log_{4}\bigg{(}&\lambda\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}|\langle\tilde{0}|^{\otimes N}\hat{\varrho}_{1}|\tilde{1}\rangle^{\otimes N}|\,\,\,+\\ &+(1-\lambda)\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}|\langle\tilde{0}|^{\otimes N}\hat{\varrho}_{2}|\tilde{1}\rangle^{\otimes N}|\bigg{)}.\end{split}$ (32)
@@
Appendix B SLOCC entanglement classes
The two pure multipartite quantum states $|\psi\rangle$ and $|\phi\rangle$ belong to the same SLOCC (stochastic local operations and classical communication) entanglement class if and only if it is possible to convert $|\psi\rangle\mapsto|\phi\rangle$ and $|\phi\rangle\mapsto|\psi\rangle$ with non-zero probabilities using local operations and classical communication, i.e., if there exist a set of invertible operators $\hat{O}_{k}$ (so that the following condition is symmetric) such that $|\phi\rangle=\bigotimes_{k=1}^{N}\hat{O}_{k}|\psi\rangle$, where $N$ is the number of constituents.
The number of entanglement classes grows rapidly with the number of parties and the local dimension. For pure states of two qubits, there are only two entanglement classes, with a Bell state as an extremal representative of one and separable states forming the other. Increasing the number of qubits to three yields two disjoint genuinely multipartite entanglement classes, together forming six different classes. Interestingly, the growth of the number of classes is not even exponential; already for four qubits, there are infinitely many entanglement classes, which can be divided into nine groups [Verstraete_2002].
Appendix C Tensor rank and multipartite entanglement classification
@@
Appendix D Entanglement classes for four qubits
@@
Appendix E Absolutely Maximally Entangled states AME(5,2) and AME(6,2)

### v2

@@ III.2 Four-qubit entanglement classification
For each of the nine classes, we maximize the correlator $\mathcal{Q}$ by picking the maximally entangled representatives of each group (Appendix C and Ref. [Vintskevich_2023]). As shown in Table 3, the highest value of the correlator is reached by the GHZ state ($|G\rangle$ group), and $|E_{1}\rangle$ and $|E_{2}\rangle$ groups, while the smallest $\mathcal{Q}$ is for group $|E_{6}\rangle$.
@@
Table 3: Value of the correlator $\mathcal{Q}$ for maximally entangled four-qubit states in nine entanglement classes (see Appendix C). The first column corresponds to the value of the correlator of the full state, while the latter four show $\mathcal{Q}$ after discarding (tracing out) one qubit. Null values corresponds to negative lack of entanglement. Observe that the state exhibiting the highest entanglement with respect to four-qubit correlator ($|G\rangle$) does not possess any three-qubit entanglement. Conversely, the states with lower four-qubit entanglement preserve high values of the three-qubit correlator. This is in agreement with similar observations concerning distribution of entanglement among different divisions [Verstraete_2002].
@@
Here, we consider only the qubit case, $d=2$. In the case of $N=2,3$ qubits, the AME state is the corresponding GHZ state. Surprisingly, there are no four-qubit AME states [Higuchi_2000], and there are no AME states for $N>6$ [Scott_2004, Huber_2017]; but they do exist for $N=5,6$ [Facchi_2008, Goyeneche_2015], taking the form of
@@
where the values of binary (±$\pm 1$) coefficients $\alpha^{(5)}_{i}$ and $\alpha^{(6)}_{i}$ are given in Appendix D. For ${\rm AME}(5,2)$ the correlator takes $\mathcal{Q}=1.7$ and for ${\rm AME}(6,2)$ it takes value $\mathcal{Q}=2.4$. These values are about half of maximal values for $N=5$ and $N=6$ qubits ($\mathcal{Q}=4$ and $\mathcal{Q}=5$ respectively, see Table 1.).
@@ Acknowledgement
We thank Guillem Müller-Rigat, Anna Sanpera, Geza Toth, and Karol Życzkowski for reading the manuscript and their useful comments. JCh was supported by the National Science Centre, Poland, within the QuantERA II Programme that has received fund ing from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733, Project No. 2021/03/Y/ST2/00195. ICFO group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA MAQS PCI2019-111828-2; QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union.
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union’s Horizon 2020 program under the Marie Sklodowska-Curie grant agreement No 847648; “La Caixa” Junior Leaders fellowships, La Caixa” Foundation (ID 100010434): CF/BQ/PR23/11980043.
@@ Appendix A Convexity of correlator $\mathcal{Q}$
$\begin{split}2\log_{4}\bigg{(}\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}&|\langle\tilde{0}|^{\otimes N}(\lambda\hat{\varrho}_{1}+(1-\lambda)\hat{\varrho}_{2})|\tilde{1}\rangle^{\otimes N}|\bigg{)}\leqslant\\ \leqslant 2\log_{4}\bigg{(}&\lambda\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}|\langle\tilde{0}|^{\otimes N}\hat{\varrho}_{1}|\tilde{1}\rangle^{\otimes N}|\,\,\,+\\ &+(1-\lambda)\max_{\{|\tilde{0}\rangle,|\tilde{1}\rangle\}}|\langle\tilde{0}|^{\otimes N}\hat{\varrho}_{2}|\tilde{1}\rangle^{\otimes N}|\bigg{)}.\end{split}$ (33)
@@
Appendix B Tensor rank and multipartite entanglement classification
@@
Appendix C Entanglement classes for four qubits
@@
Appendix D Absolutely Maximally Entangled states AME(5,2) and AME(6,2)